chlov:

the sum of the ages of thanos and his mother is 38years now after 5 years the age of thanos's mother is 3times of thanos age . Find their age now

jhonyy9:

so i dont know if i ve understood it right ,but i think it in this way : T + m = 38 38 -5 = 3*T = > T = 11

jhonyy9:

or an other version T + m = 38 and so after 5 years T +m = 38 +5 where m = 3T m - mather"s age T - Thanos"s age after 5 years T + m = 38 +5 and given m = 3T so T + 3T = 43 => 4T = 43 any idea - how we need understanding right the text of this problem ?